OpenStudy (anonymous):

Write the equation of the line that is parallel to the line 3x - y = -3 and passes through the point (4, -2).

OpenStudy (anonymous):

the line is paralell to this so has slope same as this i.e. slope=m=-3 now point slope form is.. (y-y1)=m(x-x1) =>(y--2)=-3(x-4) now solve it...yourself..

OpenStudy (anonymous):

that looks more confusing then what i asked

OpenStudy (anonymous):

oh...which thing is confusing..?@girl__in__the__shadow

OpenStudy (anonymous):

all of it i dont understand math at all

OpenStudy (anonymous):

oky..see..you have to write the eq of paralell line...and in irder to write eq of any lline you require slope and a point at least...okay..??

OpenStudy (anonymous):

yes?

OpenStudy (anonymous):

here as both eq as paralell so thier slope is same...coz whenever two lines are paralell ther slope is same...and if they are perpendicular then -ve reciprocal..okay..??